I need to know process of computing those problem and why that answer is correct.

1. An XYZ OCT 30 call option is trading at a premium of 2 and 1/2. If XYZ is trading at 28, the option has which two of the following properties?

1. An intrinsic value of 2

2. An intrinsic value of 0

3. A time value of 1/2

4. A time value of 2 and 1/2

Answer : 2 and 4

2. M. Bullock opens a margin account and purchases 100 shares of xyz at $48 per share and purchases 1 xyz 50 put at $6. if reg T is 50%, how much margin is she required to deposit?

Answer : $3,000

3. An investor purchase 1 April 60 put and 1 April 50 call. If both options finish out-of-the-money and all options are correctly priced.

a. profits will be positive

b. profits will be negative

c. profits will equal zero

Answer : d. none of the above

4. Assume that Walt Disney stock was purchased at $110 when 3 month calls were priced at $20 with an exercise price of $100. If 100 shares of stock were purchased long and 1 call option (100 shares) was shorted, the payoff at expiration if Walt Disney is priced at $114.50 is

Answer : $10,000

5. In April, a customer purchases 1 ABB Jul 85 call for 5 and purchases 1 ABB Jul 90 put for 8. ACB stock is trading at 87 if ABC stays 87 and both options are sold for intrinsic value, the customer will realize an

Answer: $500 profit

6. Your client purchases a straddle in his margin account. He pays 4 for the call and 6 for the put. what is the margin requirement for the straddle?

Answer: $1,000

7. The maximum potential loss to the following portfolio is:

Long one LMN APR 25 call at 3/4 and

Short one LMN APR 20 call at 2 and 1/4

Answer : $3.50 per share

8. If you write one XYZ Oct 20 call at 1 and 1/2 with no other underlying position, the maximum loss potential is:

Answer: Unlimited

9. Consider the following portfolio for the next question

Short one LMN Apr 25 call at 5

Short onen LMN Apr 25 put at 1

Given the above portfolio, the investor will breakeven on the above strategy if LMN trades at which two of the following?

Answer : 19 and 31
Answer #1

1) Option price has two components Intrinsic value and Time value

Intrinsic value = The amount by which the strike price is In The Money.

For a call option, Intrinsic value = max(Spot price - Strike Price, 0)

Intrinsic value = max(28 - 30, 0) = max(-2, 0) = 0

Option price = Intrinsic value + Time value

2 and 1/2 = 0 + Time value

So, Time value = 2 and 1/2.

Please do not downvote for not answering the remaining questions. As per the HOMEWORKLIB RULES, I can answer only the first question when there are many questions.

If you understand the solution, please upvote :-)